### 20 Dec 2023 v1

@@ 1 Introduction
In this framework, we attempt to provide here an additional insight into some critical aspects of disordered quantum spin-ladder systems, not fully considered previously. The goal of the present study is threefold: (i) Monitor the trend of the crossing points of the specific heat and uniform susceptibility curves, (ii) scan the response of the spin-gap parameter, and (iii) scrutinize the low-temperature behaviour of the staggered susceptibility and structure factor, while increasing the number of legs of the spin-ladder system and for different values of the disorder parameter. Whenever it is possible, we compare our results with the main outcomes of Ref. Kanbur2 .
@@ 3.1 Simulation details
Using the above scheme we generated numerical data for the quenched random-bond spin-ladder system for different system sizes at a wide temperature spectrum and for several values of the disorder strength parameter within the range $r=\{0-0.9\}$. As already mentioned above, our geometrical setup consists of $L_{x}\times L_{y}$ systems, where $L_{x}$ defines the linear dimension along the legs of the lattice and is fixed at $L_{x}=256$ during all simulations. We note in passing that we have performed some additional test simulations using $L_{x}$ values up to $512$ to identify possible finite-size and crossover effects in the observed quantities. However, our analysis, not shown here for brevity, disclosed that the results are almost $L_{x}$-independent so the considered $L_{x}=256$ dimension is enough for an accurate determination of the properties. $L_{y}$ now stands for the number of legs of the ladder system. In fact, one of the most interesting aspects of the physics of quantum Heisenberg ladders is that their low-energy properties are distinct when comparing even and odd $L_{y}$ values White . It is well-known that there is a spin gap for the case of even $L_{y}$ values, while ladder systems having odd $L_{y}$ values are gapless Sandvik1 . In this framework, we only studied systems with $L_{y}=\{2,4,6\}$. Closing, some numerical details: We performed simulations over $500$ independent random realizations for each pair of $(r,L_{y})$ values. We applied boundary conditions which are periodic along the legs and free along the rungs of the spin ladder. In our protocol, the first $10^{5}$ Monte Carlo steps were discarded during the thermalization process and numerical data were collected and analyzed during the following $5\times 10^{5}$ Monte Carlo steps. Finally, errors were estimated using the standard jackknife method Barkema ; Binder .

### 11 Mar 2024 v2

@@ 1 Introduction
In this framework, we attempt to provide here an additional insight into some critical aspects of disordered quantum spin-ladder systems, not fully considered previously. Therefore, the goal of the present study is threefold: (i) Monitor the trend of the crossing points of the specific heat and uniform susceptibility curves, (ii) scan the response of the spin-gap parameter, and (iii) scrutinize the low-temperature behaviour of the staggered susceptibility and structure factor, while increasing the number of legs of the spin-ladder system and for different values of the disorder parameter. Whenever it is possible, we compare our results with the main outcomes of Ref. Kanbur2 . Before going on further, we would like to underline that the existence of the crossing point phenomenon, observed also in the present work, is one of the most interesting and puzzling effects in strongly correlated electron systems. Up to now, and to the best of our knowledge, no unified mechanism for this phenomenon has been proposed. However, it is generally believed that crossing points occur in certain systems that are in the vicinity of a quantum or a second-order phase transition, or some physical systems showing magnetic instability, such that their properties are prominently sensitive to changes in thermodynamic variables, such as the magnetic field and the temperature.
@@ 3.1 Simulation details
Using the above scheme we generated numerical data for the quenched random-bond spin-ladder system for different system sizes at a wide temperature spectrum and for several values of the disorder strength parameter within the range $r=\{0-0.9\}$. As already mentioned above, our geometrical setup consists of $L_{x}\times L_{y}$ systems, where $L_{x}$ defines the linear dimension along the legs of the lattice and is fixed at $L_{x}=256$ during all simulations. We note in passing that we have performed some additional test simulations using $L_{x}$ values up to $512$ to identify possible finite-size and crossover effects in the observed quantities. However, our analysis, not shown here for brevity, disclosed that the results are almost $L_{x}$-independent so the considered $L_{x}=256$ dimension is enough for an accurate determination of the properties. $L_{y}$ now stands for the number of legs of the ladder system. In fact, one of the most interesting aspects of the physics of quantum Heisenberg ladders is that their low-energy properties are distinct when comparing even and odd $L_{y}$ values White . It is well-known that there is a spin gap for the case of even $L_{y}$ values, while ladder systems having odd $L_{y}$ values are gapless Sandvik1 . In this framework, we only studied systems with $L_{y}=\{2,4,6\}$. Closing, some numerical details: We performed simulations over $500$ independent random realizations for each pair of $(r,L_{y})$ values. We applied boundary conditions which are periodic along the legs and free along the rungs of the spin ladder. In our protocol, the first $10^{5}$ Monte Carlo steps were discarded during the thermalization process and numerical data were collected and analyzed during the following $5\times 10^{5}$ Monte Carlo steps. We note in passing that the simulation time needed for a single realization on a node of a Dual Intel Xeon E5-2690 V4 processor was approximately $30$ minutes for the largest system size $(L_{x}=256,L_{y}=6)$ and the minimum temperature value that we have reached during the simulations. Finally, errors were estimated using the standard jackknife method Barkema ; Binder .